# Differentially private projection-depth-based medians

### v1

@@ Abstract
We develop $(\epsilon,\delta)$-differentially private projection-depth-based medians using the propose-test-release (PTR) and exponential mechanisms. Under general conditions on the input parameters and the population measure, (e.g. we do not assume any moment bounds), we quantify the probability the test in PTR fails, as well as the cost of privacy via finite sample deviation bounds. We demonstrate our main result on the canonical projection-depth-based median. In the Gaussian setting, we show that the resulting deviation bound matches the known lower bound for private Gaussian mean estimation, up to a polynomial function of the condition number of the covariance matrix. In the Cauchy setting, we show that the “outlier error amplification” effect resulting from the heavy tails outweighs the cost of privacy. This result is then verified via numerical simulations. Additionally, we present results on general PTR mechanisms and a uniform concentration result on the projected spacings of order statistics.
@@ 1 Introduction
Early literature on differentially private median estimation focuses on the univariate setting [10, 3, 7, 24]. In the multivariate setting, several authors have studied differentially private medians through the lens of differentially private mean estimation. These works have mainly focused either on the Tukey median [6, 17, 4], or variants of it [15]. In addition, as part of a more general framework for private high-dimensional estimation, Liu et al., 2021b [18] studied one instance of a projection-depth-based median based on the trimmed mean and the trimmed variance. Given that their focus was mean estimation, all of these works enforce a bound on the number of moments possessed by the population measure. However, a classical method for assessing the robustness of a median is to examine its behavior in the case where the population measure possesses no moments. Recently, Ramsay et al., [20] considered this problem. They studied differentially private multivariate medians arising from depth functions, including the Tukey median, without enforcing any moment bounds on the population measure. However, their framework does not cover projection-depth-based medians [26, 27].
@@
Unlike the Tukey median and those studied by Ramsay et al., [20], when the location and scale measures are chosen appropriately, projection-depth-based medians simultaneously possess affine equivariance and achieve a high breakdown point [26, 27]. Affine equivariance is a critical property for location estimates, as it ensures that the units of measurement do not have an effect on the resulting estimate, see, for example, Zuo and Serfling, [28]. In addition, the breakdown point is a staple measure of the robustness of an estimator. It measures the fraction of points which must be corrupted to make an estimate arbitrarily large [14]. For instance, the breakdown point of the Tukey median degrades quickly with the dimension $(1+d)^{-1}$, whereas, in general, the projection-depth-based medians retain a much higher breakdown point, (e.g., $\approx 1/2-d/n$ [26]). One may also note that (when the location and scale measures are chosen appropriately) the projection-depth-based medians enjoy higher relative efficiency than that of the Tukey median [26]. Motivated by these properties, our present work develops and studies differentially private projection-depth-based medians. We consider a general location estimation setting and do not place any moment bounds on the population measure.
@@
Our main tool for privatizing projection-depth-based medians is the propose-test-release (PTR) algorithm, which was introduced by Dwork and Lei, [10] in the context of differentially private estimation of various robust, univariate statistics. Our work takes inspiration from that work, as well as from the works of Brunel and Avella-Medina, [7] and Brown et al., [6]. Brunel and Avella-Medina, [7] used PTR to develop sub-Gaussian median estimators in the univariate setting. Additionally, Brown et al., [6] developed a private Gaussian mean estimator, using a PTR algorithm based on safety and the exponential mechanism, both of which we incorporate into our proposed algorithm. Notably, Brown et al., ’s algorithm was also generalized by Liu et al., 2021b [18] to be an all-purpose solution to various private, statistical problems. Our work builds on that of Liu et al., 2021b [18], in the sense that their mean estimator fits into the class of projection-depth-based estimators.
@@
Our contributions are as follows: We first propose a class of $(\epsilon,\delta)$-differentially private projection-depth-based medians. Next, under general conditions on the population measure, we quantify both the probability that the PTR algorithm fails to release an estimate and the cost of privatizing the projection-depth-based medians (see Theorems 3.1, and 3.2, respectively). To measure the cost of privacy, we provide finite sample bounds on the distance between a private projection-depth-based median and the corresponding population value, which hold with high probability. We then demonstrate our methods for the canonical projection-depth-based median. In particular, we show that Theorems 3.1 and 3.2 apply provided the “projected densities” are bounded below in an interval containing a range of inner quantiles (Condition 4). As a first example, we show that in the Gaussian setting, our deviation bound matches the optimal lower bound for sub-Gaussian mean estimation given by [16, 8], up to a polynomial function of the condition number of the covariance matrix (see Example 3.1). We then show that in the Cauchy setting, the increase in error resulting from contamination or heavy tails, which we call the “outlier error amplification” is larger than the cost of privacy (see Example 3.2). We then confirm this phenomenon with numerical studies (see Section 4). Lastly, we provide some results for analysis of general PTR mechanisms (see Section 5), as well as a uniform concentration result on the projected spacings of order statistics (see Lemma C.2), which may be of general interest.
@@ 2.1 Differential privacy
We first review necessary concepts from differential privacy. For a textbook introduction to differential privacy, see Dwork and Roth, [11]. A dataset of size $n\times d$ is a collection of $n$ points in $\mathbb{R}^{d}$, $\mathbb{X}_{n}=\{X_{i}\}_{i=1}^{n}$, possibly with repetitions. Let $\mathbf{D}_{n\times d}$ be the set of all datasets of size $n\times d$. For datasets $\mathbb{X}_{n}$ and $\mathbb{Y}_{n}$, let $\mathbb{X}_{n}\triangle\mathbb{Y}_{n}$ represent the dataset containing all rows in $\mathbb{X}_{n}$ which are not contained in $\mathbb{Y}_{n}$. Define $\mathbf{A}(\mathbb{X}_{n},m)=\left\{\mathbb{Y}_{n}\in\mathbf{D}_{n\times d}\colon|\mathbb{X}_{n}\triangle\mathbb{Y}_{n}|=m\right\}$, the set of datasets of size $n\times d$ which differ from $\mathbb{X}_{n}$ in exactly $m$ rows. Define the empirical measure of $\mathbb{X}_{n}$, ${\hat{\nu}}_{\mathbb{X}_{n}}$, as that which assigns $1/n$ probability to each row in $\mathbb{X}_{n}$. For a dataset with empirical measure ${\hat{\nu}}_{\mathbb{X}_{n}}$ define
~$\widetilde{\mathcal{M}}_{1}({\hat{\nu}}_{\mathbb{X}_{n}},m)=\{\tilde{\nu}\in\mathcal{M}_{1}(\mathbb{R}^{d})\colon\tilde{\nu}={\hat{\nu}}_{\mathbb{Y}_{n}}\text{ for some }\mathbb{Y}_{n}\in\mathbf{A}(\mathbb{X}_{n},m)\},$
where $\mathcal{M}_{1}(\mathbb{R}^{d})$ is the set of measures on $\mathscr{B}(\mathbb{R}^{d})$, and $\mathscr{B}(S)$ denotes the Borel sets of $S$.
Suppose that we observe a dataset ${\mathbb{X}}_{n}$ comprised of $n$ i.i.d. samples from a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$, and wish to estimate a parameter $\theta\in\Theta$. One way to do this is via a mechanism. Formally, a mechanism, $P_{\hat{\nu}}$, maps an empirical measure $\hat{\nu}$ to $\mathcal{M}_{1}(\Theta)$. The estimator $\tilde{\theta}$ is then taken to be a draw from $P_{\hat{\nu}}$. When the corresponding measure is clear from context, we will suppress dependence on it in notation, denoting $P_{\hat{\nu}}$ as $P$. To produce a differentially private statistic, $P_{\hat{\nu}}$ should satisfy the following privacy guarantee.

### v2

@@ Abstract
We develop $(\epsilon,\delta)$-differentially private projection-depth-based medians using the propose-test-release (PTR) and exponential mechanisms. Under general conditions on the input parameters and the population measure, (e.g. we do not assume any moment bounds), we quantify the probability the test in PTR fails, as well as the cost of privacy via finite sample deviation bounds. We then present a new definition of the finite sample breakdown point which applies to a mechanism, and present a lower bound on the finite sample breakdown point of the projection-depth-based median. We demonstrate our main results on the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In the Gaussian setting, we show that the resulting deviation bound matches the known lower bound for private Gaussian mean estimation. In the Cauchy setting, we show that the “outlier error amplification” effect resulting from the heavy tails outweighs the cost of privacy. This result is then verified via numerical simulations. Additionally, we present results on general PTR mechanisms and a uniform concentration result on the projected spacings of order statistics, which may be of general interest.
@@ 1 Introduction
Early literature on differentially private median estimation focuses on the univariate setting [11, 3, 7, 25]. In the multivariate setting, several authors have studied differentially private medians through via differentially private mean estimation. These works have mainly focused either on the Tukey median [6, 17, 4], or variants of it [15]. In addition, as part of a more general framework for private high-dimensional estimation, Liu et al., 2021b [18] studied one instance of a projection-depth-based median based on the trimmed mean and the trimmed variance. Given that their focus was mean estimation, all of these works enforce a bound on the number of moments possessed by the population measure. However, a classical method for assessing the robustness of a median is to examine its behavior in the case where the population measure possesses no moments. Recently, Ramsay et al., [21] considered this problem. They studied differentially private multivariate medians arising from depth functions, including the Tukey median, without enforcing any moment bounds on the population measure. However, their framework does not cover projection-depth-based medians [27, 28].
@@
Unlike the Tukey median and those studied by Ramsay et al., [21], when the location and scale measures are chosen appropriately, projection-depth-based medians simultaneously possess affine equivariance and achieve a high breakdown point [27, 28]. Affine equivariance is a critical property for location estimators, as it ensures that the units of measurement do not have an effect on the resulting estimate, see, for example, Zuo and Serfling, [29]. In addition, the breakdown point is a common measure of the robustness of an estimator. It measures the fraction of points which must be corrupted to make an estimate arbitrarily large [14]. For instance, the breakdown point of the Tukey median degrades quickly with the dimension, $(1+d)^{-1}$, whereas the projection-depth-based medians generally retain a much higher finite sample breakdown point, (e.g., approximately $1/2-d/n$ [27]). This finite sample breakdown point is carried over to the private setting, (see Theorem 3.3). One may also note that, when the location and scale measures are chosen appropriately, the projection-depth-based medians enjoy higher relative efficiency than that of the Tukey median [27]. Motivated by these properties, our work develops and studies differentially private projection-depth-based medians. We consider a general location estimation setting and do not place any moment bounds on the population measure.
@@
Our main tool for privatizing projection-depth-based medians is the propose-test-release (PTR) algorithm, which was introduced by Dwork and Lei, [11] in the context of differentially private estimation of various robust, univariate statistics. Our work takes inspiration from that work, as well as from the works of Brunel and Avella-Medina, [7] and Brown et al., [6]. Brunel and Avella-Medina, [7] used PTR to develop sub-Gaussian median estimators in the univariate setting, and Brown et al., [6] developed a private Gaussian mean estimator, using a PTR mechanism based on safety and the exponential mechanism. We incorporate both ideas into our proposed mechanism. Notably, the mechanism proposed by Brown et al., [6] was generalized by Liu et al., 2021b [18] to be an all-purpose solution to various private, statistical problems. Our work builds on that of Liu et al., 2021b [18], in the sense that their mean estimator fits into the class of projection-depth-based estimators.
@@
Our contributions are as follows: We first propose a class of $(\epsilon,\delta)$-differentially private projection-depth-based medians. Next, under general conditions on the population measure, we quantify both the probability that the PTR mechanism fails to release an estimate and the cost of privatizing the projection-depth-based medians (Theorems 3.1, and 3.2, respectively). To measure the cost of privacy, we provide finite sample bounds on the distance between a private projection-depth-based median and the corresponding population value, which hold with high probability. Next, we present a new definition of the finite sample breakdown point which applies to a mechanism, and present a lower bound on the finite sample breakdown point of the projection-depth-based median (Theorem 3.3). We then demonstrate our methods for the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In particular, we show that Theorems 3.1 and 3.2 apply provided the “projected densities” are bounded below in an interval containing a range of inner quantiles (Condition 4). As a first example, we show that in the Gaussian setting, our deviation bound matches the optimal lower bound for sub-Gaussian mean estimation given by Kamath et al., [16] and Cai et al., [8] (Example 3.1). We then show that in the Cauchy setting, the increase in error resulting from contamination or heavy tails, which we call the “outlier error amplification” is larger than the cost of privacy (Example 3.2). We confirm this phenomenon with numerical studies (Section 4). Lastly, we provide some results for analysis of general PTR mechanisms (Section 5), as well as a uniform concentration result on the projected spacings of order statistics (Lemma C.2), which may be of general interest.
@@ 2 Preliminaries
2.1 Outlyingness and projection-depth-based medians
We first review projection-depth-based medians, starting with projected outlyingness [9, 29, 27]. Let $\mathcal{M}_{1}(\mathbb{R}^{d})$ be the set of measures on $\mathscr{B}(\mathbb{R}^{d})$, where $\mathscr{B}(S)$ denotes the Borel sets of $S$. For a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and $u$ from the unit $(d-1)$-sphere, $\mathbb{S}^{d-1}$, let $\nu_{u}$ be the law of $X^{\top}u$ if $X\sim\nu$.
Definition 2.1.
[…]
The projection-depth-based medians enjoy several nice properties, including local and global robustness, affine equivariance, strong $\sqrt{n}$-consistency, and a relatively simple limiting distribution [27, 28]. To the best of our knowledge, among the competing affine equivariant medians, when the location and scale measures are chosen appropriately, (e.g. the median and median absolute deviation), the projection-depth-based medians have the highest breakdown point [27]. We show later that the private projection-depth-based medians also admit high finite sample breakdown points, (Theorem 3.3). Projection-depth-based medians have been shown to have higher asymptotic efficiency, relative to competitors [27]. For more information on projection-depth-based medians, we refer the reader to [27, 28].
@@ 2.2 Differential privacy
Next, we review necessary concepts from differential privacy. For a textbook introduction to differential privacy, see Dwork and Roth, [12]. A dataset of size $n\times d$ is a collection of $n$ points in $\mathbb{R}^{d}$, $\mathbb{X}_{n}=\{X_{i}\}_{i=1}^{n}$, possibly with repetitions. Let $\mathbf{D}_{n\times d}$ be the set of all datasets of size $n\times d$. For datasets $\mathbb{X}_{n}$ and $\mathbb{Y}_{n}$, let $\mathbb{X}_{n}\triangle\mathbb{Y}_{n}$ represent the dataset containing all rows in $\mathbb{X}_{n}$ which are not contained in $\mathbb{Y}_{n}$. Define $\mathbf{S}(\mathbb{X}_{n},m)=\left\{\mathbb{Y}_{n}\in\mathbf{D}_{n\times d}\colon|\mathbb{X}_{n}\triangle\mathbb{Y}_{n}|=m\right\}$, the set of datasets of size $n\times d$ which differ from $\mathbb{X}_{n}$ in exactly $m$ rows. Define the empirical measure of $\mathbb{X}_{n}$, ${\hat{\nu}}_{\mathbb{X}_{n}}$, as that which assigns $1/n$ probability to each row in $\mathbb{X}_{n}$. For a dataset with empirical measure ${\hat{\nu}}_{\mathbb{X}_{n}}$ define
~$\widetilde{\mathcal{M}}_{1}({\hat{\nu}}_{\mathbb{X}_{n}},m)=\{\tilde{\nu}\in\mathcal{M}_{1}(\mathbb{R}^{d})\colon\tilde{\nu}={\hat{\nu}}_{\mathbb{Y}_{n}}\text{ for some }\mathbb{Y}_{n}\in\mathbf{S}(\mathbb{X}_{n},m)\}.$
Suppose that we observe a dataset ${\mathbb{X}}_{n}$ comprised of $n$ i.i.d. samples from a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$, and wish to estimate a parameter $\theta\in\Theta$. One way to do this is via a mechanism. Formally, a mechanism, $P_{\hat{\nu}}$, maps an empirical measure $\hat{\nu}$ to $\mathcal{M}_{1}(\Theta)$. The estimator $\tilde{\theta}$ is then taken to be a draw from $P_{\hat{\nu}}$. When the corresponding measure is clear from context, we will suppress it in the notation, denoting $P_{\hat{\nu}}$ as $P$. To produce a differentially private statistic, $P_{\hat{\nu}}$ should satisfy the following privacy guarantee.